Eureka is 12 years younger than Renton. In 3 years Renton will be four times as old as Eureka. How old are they now?
Answer by lwsshak3(11628)

Eureka is 12 years younger than Renton. In 3 years Renton will be four times as old as Eureka. How old are they now?
***
let x=Eureka's age
x+12=Renton's age
..
In 3 yrs:
x+3=Eureka's age
x+12+3=x+15=Rentom's age
x+15=4(x+3)
x+15=4x+12
3x=3
x=1
x+12=13
Eureka's age=1
Renton's age=13